OpenStudy (anonymous):

Write the trigonometric expression in terms of sine and cosine, and then simplify. tan(theta) // sec(theta) - cos(theta)

OpenStudy (anonymous):

tan = sin/cos sec = 1/cos

OpenStudy (amistre64):

nite lol...and you too star :) librarys closing up on me...

OpenStudy (anonymous):

Are you literally in a library?

OpenStudy (anonymous):

Thanks a lot amistre :)

OpenStudy (anonymous):

You can plug in the given suzi has given you and you'll get the following:\[\tan \theta /(\sec \theta - \cos \theta) = (\sin \theta/\cos \theta)/[(1-\cos^2 \theta)/\cos \theta]\] now simplify : \[= (\cos \theta \sin \theta)/(\cos \theta \sin^2 \theta)\] \[= 1/\sin \theta\] ^_^

OpenStudy (anonymous):

good night amistre :)